OpenStudy (anonymous):

√5+x>3 is this alwasys,sometimes or never true

OpenStudy (anonymous):

Can you rearrange the equation to make \(x\) the subject?

OpenStudy (anonymous):

Identify if each statement is always true,sometimes true or never true over the real numbers for which it is defined √5+x>3

OpenStudy (anonymous):

So rearrange the equation to get \(x>3-\sqrt{5}\) (by taking \(\sqrt{5}\) from both sides). It's still the same equation, just rearranged. Can you think of a value of \(x\) that makes this true? How about \(5-\sqrt{5}\)? This means that it's (at least) sometimes true. Let's see if there's values of \(x\) for which this inequality isn't true. How about \(x= 1-\sqrt{5}\)? So we can see that it is sometimes true, and sometimes not true, depending on the value of \(x\) taken.